OpenStudy (anonymous):

a culture contains 1500 bacteria initially and doubles every 30 mins. a) find a function that models the number of bacteria at time t. b) find the number of bacteria after 2 hours. c) after how much time will there be 4000 bacteria

OpenStudy (diana):

f(x)= 1500 + 30t f(4)= 1500+30(4)=1620 4000=1500+30t= 250/3

OpenStudy (anonymous):

That's not correct

OpenStudy (diana):

Why?

OpenStudy (deoxna):

f(x)=1500*2^(2t) where t is a value in hours and 30 min=0.5 hours f(2)=1500*16=24000 4000=1500*2^(2t) 8/3=2^(2t) ln 8/3=2t ln 2 (ln 8/3)/(2 ln 2)= t t=0.708

OpenStudy (anonymous):

That's if you have 1500 bacteria and you have another 30 bacteria every hour

OpenStudy (anonymous):

DNA's answer is correct

OpenStudy (diana):

oh i see